Simplify

5/(x-2) + 1/ 4x - 8



1
Expert's answer
2021-03-05T01:16:52-0500

"\frac{5}{x-2}+\frac{1}{4x-8}\:"


"\frac{5\left(4x-8\right)+\left(x-2\right)}{\left(x-2\right)\left(4x-8\right)}"


"\frac{\left(20x-40\right)+\left(x-2\right)}{\left(x-2\right)\left(4x-8\right)}"


"\frac{20x-40+x-2}{\left(x-2\right)\left(4x-8\right)}\:"


"\frac{20x+x-40-2}{x\left(4x-8\right)-2\left(4x-8\right)}"


"\frac{21x-42}{4x^2-8x-8x+16}"


"\frac{21x-42}{4x^2-16x+16}"


"\frac{21x-42}{4\left(x-2\right)\left(x-2\right)}"


"\frac{21\left(x-2\right)}{4\left(x-2\right)\left(x-2\right)}"


"\frac{21}{4\left(x-2\right)}"


"\frac{21}{4x-8}"